GNU gv Manual

First Edition, Updated for gv Version 3.6.6. Johannes Plass José E. Marchesi This is the First edition of the GNU gv Manual, updated for gv version 3.6.6. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995, 1996, 1997 Johannes Plass Copyright \begin{matrix}c \\ ◯ \end{matrix} 2004 José E. Marchesi

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU General Public License, Version 3 or any later version published by the Free Software Foundation.

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Invoking gv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Resources. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3.1 Resource files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Resources of gv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4 Mouse and key bindings . . . . . . . . . . . . . . . . . . . . . . 12

4.1 Mouse bindings in the Main and the Zoom window . . . . . . . . . . . . 12 4.2 Mouse bindings in the File Selection popup . . . . . . . . . . . . . . . . . . . . 12 4.3 Mouse bindings in the Table of Contents . . . . . . . . . . . . . . . . . . . . . . 13 4.4 Mouse bindings in the Panner widget . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.5 Mouse bindings in the << and >> button windows. . . . . . . . . . . . . . 13 4.6 Key Bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

4.6.1 Notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.6.2 Miscellaneous keys:. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.6.3 Orientation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.6.4 Magnification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.6.5 Navigating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

5 Comments about the user interface . . . . . . . . . . 17

5.1 Scales. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.2 Scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.3 Displaying new versions of a document . . . . . . . . . . . . . . . . . . . . . . . . 18

6 Frequently asked questions . . . . . . . . . . . . . . . . . . . 19

6.1 What happened to the scroll bars? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.2 How can I make even more room for the image? . . . . . . . . . . . . . . . 19 6.3 Small characters are not very clear . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.4 The output is not refreshed automatically . . . . . . . . . . . . . . . . . . . . . 19

7 gv internals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

7.1 Interface with ghostscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 7.1.1 Ghostview interface to ghostscript . . . . . . . . . . . . . . . . . . . . . . . . 20 7.1.2 Drawing on a Window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 7.1.3 Drawing on a Pixmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 7.1.4 The GHOSTVIEW environment variable . . . . . . . . . . . . . . . . . 20 7.1.5 The GHOSTVIEW property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 7.1.6 Events from ghostscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

7.2 Comments recognized by ghostview. . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7.3 Paper Keywords and paper size in points . . . . . . . . . . . . . . . . . . . . . . 23

ii

Appendix A GNU General Public License …. 24

Chapter 1: Introduction 1

1 Introduction

GNU gv allows to view and navigate through PostScript and PDF documents on an X display by providing a user interface for the ghostscript interpreter.

Please note that gv is derived from Tim Theisen’s ghostview.

2 Invoking gv

Usage: gv [OPTION]... [FILE]
PostScript and PDF viewer.
[-]-monochrome                      display document using only black and white
[-]-grayscale                       display document without colors
[-]-color                           display document as usual
[-]-safer                           start ghostscript in safe mode
[-]-nosafer                         do not start ghostscript in safe mode
[-]-quiet                           start ghostscript with the -dQUIET option
[-]-noquiet                         do not start ghostscript with the -dQUIET option
[-]-infoSilent                      do not show any messages in the info popup window
[-]-infoErrors                      do not show warning messages in the info popup window
[-]-infoAll                         do show all messages in the info popup window
[-]-arguments=ARGS                  start ghostscript with additional options as specified
                                    by the string ARGS
[-]-page=LABEL                      display the page with label LABEL first
[-]-center                          the page should be centered automatically
[-]-nocenter                        the page should not be centered automatically
[-]-media=MEDIA                     selects the paper size to be used
[-]-orientation=ORIENTATION         sets the orientation of the page
[-]-scale=N|f.f                     selects the scale N, or arbitrary scale f.f
[-]-scalebase=N                     selects the scale base N
[-]-swap                            interchange the meaning of the orientations landscape
                                    and seascape
[-]-noswap                          do not interchange the meaning of the orientation
                                    landscape and seascape
[-]-antialias                       use antialiasing
[-]-noantialias                     do not use antialiasing
[-]-dsc                             dsc comments are respected
[-]-nodsc                           dsc comments are not respected
[-]-eof                             ignore the postscript EOF comment while scanning
                                    documents
[-]-noeof                           do not ignore the postscript EOF comment while
                                    scanning documents
[-]-pixmap                          use backing pixmap
[-]-nopixmap                        do not use backing pixmap
[-]-watch                           watch the document file for changes
[-]-nowatch                         do not watch the document file for changes
[-]-help                            print a help message and exit
[-]-usage                           print a usage message and exit
[-]-resize                          fit the size of the window to the size of the page
[-]-noresize                        do not fit the size of the window to the size of the page
-geometry [<width>][x<height>][{+-}<xoffset>{+-}<yoffset>]
[-]-ad=FILE                         read and use additional resources from FILE
[-]-style=FILE                      read and use additional resources from FILE. These resources
                                    have lower priority than those provided on the context of --ad

[-]-password=PASSWORD               Sets the password for opening encrypted PDF files
[-]-spartan                         shortcut for --style=gv_spartan.dat
[-]-widgetless                      shortcut for --style=gv_widgetless.dat
[-]-fullscreen                      start in fullscreen mode (needs support from WM)
[-]-presentation                    Presentation mode (fullscreen, Fit to window,
                            widgetless and no resizing of window)
[-]-version                         show gv version and exit

file The name of the file to be displayed. The file ending ‘.ps’ may be omitted. file may point to a compressed file (gzipped, bzipped, zipped or compressed). Viewing PDF files requires at least ghostscript version 4.x.

--ad=file

Read and use additional resources from the file file. These resources have higher priority than those provided in the context of the --style option.

--antialias, --noantialias

Whether to use antialiasing.

--arguments=args

Start ghostscrip with additional options as specified by the string args.

--center, --nocenter

Whether the page should be centered automatically.

--dsc, --nodsc

Determines if document structuring convention (DSC) comments should be respected or ignored. If --nodsc is used gv will not attempt to examine the structure of the document but will pass the file to the ghostscript interpreter as a whole. In this case no page numbers are shown and freely moving around in the document is not possible. This option may help when viewing files not conforming to the document structuring conventions.

--eof, --noeof

Defines the behaviour of the postscript scanner used to examine the structure of the document. If --noeof is used the scanner will ignore end of file (EOF) comments. This may help when viewing documents which import other documents without enclosing them within the proper "BeginDocument" and "EndDocument" comments. If --eof is used, the scanner treats an EOF comment as marking the end of the file.

--fullscreen

Selects the fullscreen mode.

--infoSilent

Do not show any ghostscript messages in an info popup window.

--infoErrors

Only show error messages from ghostscript in an info popup window.

--infoAll

Shows all ghostscript messages in an info popup window.

--pixmap, --nopixmap

If --pixmap is used gv tries to maintain off-screen regions of the displayed page by allocating a sufficiently large pixmap. If --nopixmap is used the X Server is responsible for maintaining obscured portions of the displayed page (see also the useBackingPixmap resource).

--version

Print the version number of gv to standard output and exit.

--help, --usage

Display a short help message and exit.

--scale=n

n must be an integer or a positive float. In case n is an integer, n selects the scale entry n relative to the scale 1.0. With the special value of -1000 Fit to page is selected, and likewise with the value of -1001 Fit widthto page is selected. In case n is a float, n selects an arbitrary scale.

--scalebase n

Selects the scale base n.

--monochrome, --grayscale, --color

Sets the color palette to be used.

--media=media

Selects the paper size to be used. Valid values are names of paper sizes that appear in the list of page medias as given by the medias resource.

--page=label

Display the page with label label first.

--orientation=orientation

Sets the orientation. Valid values for orientation are portrait, landscape,

seascape and upsidedown.
--quiet, --noquiet

Whether to start ghostscript with the -dQUIET option.

--resize, --noresize

Whether ghostscrip is allowed to automatically fit the size of its window to the size of the displayed page.

--password=PASSWORD

When opening a file, use PASSWORD to decrypt it. This Option is usefull for viewing encrypted PDF files.

--presentation
Is actually a shortcut for --fullscreen --noresize --scale=-1000
--widgetless.
--safer, --nosafer

Whether to start ghostscript with the -dSAFER option.

--spartan
This is a shortcut for --style=gv_spartan.dat.

--style=file

Read and use additional resources from the file file. These resources have lower priority than those provided in the context of the --ad option.

--swap, --noswap

Whether to interchange the meaning of the orientations landscape and seascape.

--watch, --nowatch

The --watch option causes gv to check the document periodically. If changes are detected gv will automatically display the newer version of the file. The document is by default checked once every second. This can be changed via the watchFileFrequency resource. Note that sending gv the SIGHUP signal also causes an update of the displayed document. This method may be used by document creators to trigger gv remotely.

--widgetless
This is a shortcut for --style=gv_widgetless.dat.

3 Resources

3.1 Resource files

In general gv does not depend on any external resource files. However, when starting gv, preferences are read from

a system specific resource file.

The name of this file is GV and is installed on as part of gv’s architecture independent library. Usually the library files are installed on ‘/usr/local/lib/gv’. Administrators may want to modify this file in order to set up gv according to the local needs of their system.

a user specific resource file.

The name of this file is either ‘~/.gv’ or as described by the XUSERFILESEARCHPATH environment variable.

The files ‘gv_user.ad’ and ‘gv_system.ad’ (located in the library directory of gv, which is most probably either ‘/usr/local/lib/gv/’ or ‘/usr/lib/gv/’) may serve as a basis for constructing the user and system specific resource files.

Note that the user and system specific resource files are not the only sources of preferences taken into account when gv is started. However, in practice these are the most important.

3.2 Resources of gv

The following describes some of the resources of gv. The precise syntax of some of the resource values may be inferred from the appended default system specific resource file.

antialias Whether antialiasing should be used. Valid values are True and False. It defaults to True.

autoCenter Whether the displayed page should automatically be centered (for instance when opening a new file). Valid values are True and False. It defaults to True.

autoResize Whether gv is allowed to resize the main window according to the size of the current page. Valid values are True and False. It defaults to True.

confirmPrint Whether printing requires an extra confirmation. Valid values are True and

False.

It defaults to True.

confirmQuit Whether leaving gv requires extra confirmation. Valid values are 0 (Never), 1 (When processing) and 2 (Always). The default value is 1, which causes a confirmation request when trying to leave gv in the presence of pending PDF to Postscript conversions. scrollingEyeGuide

If this resource is True scrolling via the keyboard causes a temporary border to be drawn around the previously visible area of the page. Allowed values are

True and False.

It defaults to True.

infoVerbose Defines the behavior of the info popup which shows messages from ghostscript. Valid values are Silent (Never), Errors (show only errror messages) and All (Any message is being shown). It defaults to Errors.

xinerama Defines whether the display resolution is determined by xinerama. Valid values are On (use Xinerama if available), Off (do not use Xinerama) and Auto (heuristically decide whether to use Xinerama or not). It defaults to Auto.

ignoreEOF Defines the behaviour of the postscript scanner as described in the Options section. Valid values are True and False. It defaults to True.

respectDSC Determines whether gv should attempt to respect document structuring comments. Valid values are True and False. It defaults to True.

reverseScrolling Defines the interpretation of directions when scrolling with the mouse. Valid values are True and False. It defaults to False.

swapLandscape Whether the meaning of landscape and seacape should be interchanged. Valid values are True and False. It defaults to False.

scratchDir Specifies a directory used to store temporary data.

It defaults to the invocation directory.

defaultSaveDir Specifies the default destination directory for files to be saved. It defaults to the home directory, ‘~/’.

useBackingPixmap If the value of this resource is False the X Server is advised that saving offscreen pixels of the displayed page would be beneficial (maintaining backing store). In this case gv does not actively maintain the contents of the page but relies on the X Server. The server is, however, always free to stop maintaining backing backing store. If the value of this resource is True the X Server is advised that maintaining backing store is not useful. In this case gv attempts to allocate a sufficiently large pixmap to store the contents of the displayed page. It defaults to True.

watchFile Whether the document should be checked periodically and updated automatically if changes are detected. Valid values are True and False. It defaults to False.

watchFileFrequency Defines the time in milliseconds elapsing between successive checks of the document when watchFile is set to True. Allowed values must be larger than 500. It defaults to 1000.

printCommand The command used for printing a file. The command may contain multiple %s wildcards which will be replaced by the name of the file to be printed. It defaults to lpr.

saveposFilename This file is used for saving positions by the z and u keys. It defaults to ~/gv.savepos.

gsInterpreter The command used to start the ghostscript interpreter. It defaults to gv.

gsCmdScanPDF The command used to extract document structuring comments from a PDF file.

It defaults to gs -dNODISPLAY -dQUIET -sPDFname=%s -sDSCname=%s
pdf2dsc.ps -c quit

gsCmdConvPDF The command used to convert a PDF file to PostScript.

It defaults to gs -dNODISPLAY -dQUIET -dNOPAUSE -sPSFile=%s %s -c quit

gsX11Device The ghostscript command line option which activates the X11 device.

It defaults to -sDEVICE=x11

gsX11AlphaDevice The ghostscript command line option which activates the X11 device with antialiasing enabled.

It defaults to -dNOPLATFONTS -sDEVICE=x11alpha

gsSafer Determines whether ghostscript should be started with the -dSAFER command line option. It defaults to True. gsQuiet Determines whether ghostscript should be started with the -dQUIET command

line option. It defaults to True.

gsArguments May be used to provide additional command line options for the ghostscript interpreter. Starting gv with the -arguments arguments option will override this resource setting. It defaults to the empty string.

magMenu Describes the menu that pops up when preparing to to zoom an area of the displayed page.

dirs Defines a list of directories accessible via the Directories menu in the file selection window. The special value Home corresponds to the users home directory, the value Tmp corresponds to the scratch directory as defined by the scratchDir resource.

It defaults to Home Tmp /usr/doc /usr/local/doc

filter Defines the default filter to be used when displaying a directory contents in the

file selection window. A filter is specified according to the syntax

   <filter>    := [<filespecs>] [no <filespecs>]
   <filespecs> := <filespec> [<filespecs>]
   <filespec>  := filename possibly including wildcards '*'
                  which match any character.

Example: The filter

   GV*filter: *.ps *.pdf no .*

screens out all files with names starting with a dot and keeps of the remaining ones only those which end on ‘.ps’ or ‘.pdf’. It defaults to no .*

filters Defines a list of filters offered in the Filters menu of the file selection window. The filter None has a special meaning causing all files to be displayed when the corresponding menu entry is selected.

miscMenuEntries Describes the menu that pops up when clicking with the third mouse button on the displayed page. The value of this resource is a list of menu entries. Allowed entries are

  • update
  • reload
  • toggle_current
  • toggle_even
  • toggle_odd
  • unmark
  • stop
  • print_all

  • print_marked
  • save_all
  • save_marked
  • line

titleStyle Whether the name of the displayed document should be shown in the title bar of the window. The name of gv’s icon will also change accordingly if this resource is set to True. If this is 0, no title will be shown. 1 means to show a the document title or the file name if there is no document title. Only the file name is shown with 2. It defaults to 1.

maximumWidth, maximumHeight When resizing gv will not attempt to exceed the size specified by these resources. These resources may be specified as a positive integer or as screen, optionally combined with a positive or negative offset. The default values listed above provide examples for this syntax. screen will automatically be replaced by the size of the screen.

maximumWidth defaults to screen-20. maximumHeight defaults to screen-44.

minimumWidth, minimumHeight Defines a minimum size for the main window. Valid values for both resources are positive integers larger than 200.

minimumWidth defaults to 400. minimumHeight defaults to 430.

scale Sets the initial scale. The value of this resource is the offset of the scale to be selected relative to the scale 1.0 in the the Scale menu. It defaults to 0.

scaleBase Selects the initial scale base. The value of this resource should be a positive integer. It defaults to 1.

scales Defines the available scales bases and scales in the Scale menu. orientation

Sets the initial orientation. Valid values are

  • portrait
  • landscape
  • seascape
  • upside-down
  • automatic

automatic causes gv to attempt to derive the correct orientation from document structuring comments. It defaults to automatic. fallbackOrientation

Defines the paper-size to be used when automatic orientation detetction fails.

Valid values are portrait, landscape, seascape and upside-down.

It defaults to portrait.

medias A list describing the page medias known to gv. List entries starting with ! or # will not appear in the Media menu but will still be used for automatic paper size detection.

pagemedia Sets the paper size. Allowd values are as given in the above list of paper-sizes. Specifying automatic causes gv to attempt to derive the correct paper-size from document structuring comments. It defaults to automatic.

fallbackPageMedia Defines the paper-size to be used when automatic paper-size detetction fails. Valid values are as given in the above list of paper-sizes. It defaults to a4.

4 Mouse and key bindings

4.1 Mouse bindings in the Main and the Zoom window

The following mouse events are defined when the mouse pointer is either on the displayed page or on a zoomed area:

press button 1, move mouse, release button 1 Scrolls the displayed page ’proportionally’.

press button 3, move mouse, release button 3 Scrolls the displayed page ’absolutely’.

double-click with button 1 In the main window this displays the previous/next page if the double-click occured in the left/right half of the window. In a zoom window it does nothing.

press button 2, release button 2 Pops up a small menu which allows to choose a magnification for a certain area around the current mouse position. After selecting a magnification a zoom window pops up displaying the area at the chosen scale.

press button 2, move mouse, release button 2 This draws and thereby defines a rectangular region which can be displayed in a zoom window. The magnification for the area can be selected by means of a popup menu which appears after releasing button 2.

press button 1, press button 2 Reload the current page. This event sequence works only in the main window.

press button 3, release button 3 Pops up a small menu which offers a few standard actions like ’Reload’, ’Mark Page’, etc.

4.2 Mouse bindings in the File Selection popup

The following mouse events are defined when the mouse pointer is in the window of either a file or a directory list:

press button 1, move mouse, release button 1 Scrolls a file or directory list ’proportionally’.

press button 2, move mouse, release button 2 Scrolls a file or directory list ’absolutely’.

press button 2, release button 2 Scrolls a file or directory list one page up/down if the button was pressed in the upper/lower half of the list.

double-click with button 1 on a filename Selects the file and closes the File Selection popup.

click with button 1 on a directory name Causes the contents of this directory to be displayed.

4.3 Mouse bindings in the Table of Contents

The following mouse events are defined when the mouse pointer is in the window showing a list of page numbers of the current file (table of contents):

press button 1, move mouse, release button 1 Scrolls the table of contents.

press button 1 on a page number Go to this page.

press button 2 on a page number Mark this page if it is unmarked, but unmark it if it is marked.

press button 2, move mouse, release button 2 Marks/unmarks all unmarked/marked pages in the region crossed by the mouse pointer during the movement.

4.4 Mouse bindings in the Panner widget

The Panner widget is the rectangular region located close to the left edge of the main window. It indicates the size and the position of the visible area of the current page relative to the total page.

The following mouse events are defined when the mouse pointer is in this region:

press button 1, move mouse Scrolls the displayed page.

press button 2 Display the previous page.

press button 3 Display the next page.

4.5 Mouse bindings in the << and >> button windows

The << and >> buttons are used to move to another page. The following mouse events are defined:

press button 1 Move 1 page forward/backwards.

press button 2 Move 5 pages forward/backwards.

press button 3 Move 10 pages forward/backwards.

4.6 Key Bindings

The following key events are defined in the main window. Those bindings scrolling the page are also defined in zoom popups.

4.6.1 Notation

⟨\underline{\overline{S-X}}^{⟩} means press ^{⟨}\underline{\overline{Shift}}^{⟩} and key ^{⟨}\underline{\overline{X}}^{⟩} ⟨\underline{\overline{C-X}}^{⟩} means press ^{⟨}\underline{\overline{Ctrl}}^{⟩} and key ^{⟨}\underline{\overline{X}}^{⟩} ⟨\underline{\overline{SC-X}}^{⟩} means press ^{⟨}\underline{\overline{Shift}}^{⟩} or ^{⟨}\underline{\overline{Ctrl}}^{⟩} and key ^{⟨}\underline{\overline{X}}^{⟩}

4.6.2 Miscellaneous keys:

⟨\underline{\overline{A}}^{⟩} Toggle antialiasing on and off ⟨\underline{\overline{O}}^{⟩} Open a new file ⟨\underline{\overline{Q}}^{⟩} Quit gv ⟨\underline{\overline{R}}^{⟩} Toggle gv’s resizing behaviour ⟨\underline{\overline{I}}^{⟩} Respect/Ignore document structuring ⟨\underline{\overline{W}}^{⟩} Watch file / Don’t watch file ⟨\underline{\overline{S}}^{⟩} Save the marked pages ⟨\underline{\overline{SC-S}}^{⟩} Save the current file ⟨\underline{\overline{P}}^{⟩} Print the marked pages ⟨\underline{\overline{SC-P}}^{⟩} Print the current file ⟨\underline{\overline{c-L}}^{⟩} Reload the current page ⟨\underline{\overline{.}}^{⟩} Reload the current page ⟨\underline{\overline{SC-.}}^{⟩} Reopen the current file ⟨\underline{\overline{M}}^{⟩} Mark the current page ⟨\underline{\overline{N}}^{⟩} Unmark the current page ⟨\underline{\overline{Z}}^{⟩} append the current coordinate to a file whose name is given by the

          GV.saveposFilename resource.

⟨\underline{\overline{U}}^{⟩} enter a command with the current coordinate as its default parameter. By defalt it is % TODO \PutAtPos(x,y)(), where x,y are the current coordinate and the user can put needed commands into (). This command is appended to a file whose name is given by the GV.saveposFilename resource.

4.6.3 Orientation

⟨\underline{\overline{7}}^{⟩} Orientation portrait ⟨\underline{\overline{8}}^{⟩} Orientation landscape ⟨\underline{\overline{9}}^{⟩} Orientation upside-down ⟨\underline{\overline{0}}^{⟩} Orientation seascape

4.6.4 Magnification

⟨\underline{\overline{SC-6}}^{⟩} Select the scale entry -6 relative to the scale 1.0 ⟨\underline{\overline{SC-5}}^{⟩} Select the scale entry -5 relative to the scale 1.0 ⟨\underline{\overline{SC-4}}^{⟩} Select the scale entry -4 relative to the scale 1.0 ⟨\underline{\overline{SC-3}}^{⟩} Select the scale entry -3 relative to the scale 1.0 ⟨\underline{\overline{SC-2}}^{⟩} Select the scale entry -2 relative to the scale 1.0 ⟨\underline{\overline{SC-1}}^{⟩} Select the scale entry -1 relative to the scale 1.0 ⟨\underline{\overline{^}}^{⟩} Select the scale 1.0 (for german keyboard layout) ⟨\underline{\overline{‘}}^{⟩} Select the scale 1.0 ⟨\underline{\overline{1}}^{⟩} Select the scale entry 1 relative to the scale 1.0 ⟨\underline{\overline{2}}^{⟩} Select the scale entry 2 relative to the scale 1.0 ⟨\underline{\overline{3}}^{⟩} Select the scale entry 3 relative to the scale 1.0 ⟨\underline{\overline{4}}^{⟩} Select the scale entry 4 relative to the scale 1.0 ⟨\underline{\overline{5}}^{⟩} Select the scale entry 5 relative to the scale 1.0 ⟨\underline{\overline{6}}^{⟩} Select the scale entry 6 relative to the scale 1.0 ⟨\underline{\overline{+}}^{⟩} Select the next scale entry ⟨\underline{\overline{=}}^{⟩} Select the next scale entry ⟨\underline{\overline{-}}^{⟩} Select the previous scale entry

4.6.5 Navigating

⟨\underline{\overline{Arrow}}^{⟩} Scroll in the direction of the arrow ⟨\underline{\overline{S-Up}}^{⟩} Scroll columns up or jump -1 pages ⟨\underline{\overline{S-Down}}^{⟩} Scroll columns down or jump 1 pages ⟨\underline{\overline{S-Left}}^{⟩} Scroll rows left or jump -1 pages ⟨\underline{\overline{S-Right}}^{⟩} Scroll rows right or jump 1 pages ⟨\underline{\overline{C-Up}}^{⟩} Jump -1 pages (to top/left corner) ⟨\underline{\overline{C-Down}}^{⟩} Jump 1 pages (to top/left corner) ⟨\underline{\overline{C-Left}}^{⟩} Jump -5 pages (to top/left corner) ⟨\underline{\overline{C-Right}}^{⟩} Jump 5 pages (to top/left corner) ⟨\underline{\overline{D}}^{⟩} Identical to arrow up ⟨\underline{\overline{X}}^{⟩} Identical to arrow down ⟨\underline{\overline{Z}}^{⟩} Identical to arrow left ⟨\underline{\overline{Y}}^{⟩} Identical to arrow left ⟨\underline{\overline{C}}^{⟩} Identical to arrow right ⟨\underline{\overline{V}}^{⟩} Center the page ⟨\underline{\overline{SC-Space}}^{⟩} Scroll columns up or jump -1 pages ⟨\underline{\overline{Space}}^{⟩} Scroll columns down or jump 1 pages ⟨\underline{\overline{BackSpace}}^{⟩} Scroll columns up or jump -1 pages ⟨\underline{\overline{Insert}}^{⟩} Jump -5 pages ⟨\underline{\overline{Delete}}^{⟩} Jump 5 pages ⟨\underline{\overline{Home}}^{⟩} Go to the first page ⟨\underline{\overline{End}}^{⟩} Go to the last page ⟨\underline{\overline{Prior}}^{⟩} Jump -1 pages ⟨\underline{\overline{Next}}^{⟩} Jump 1 pages ⟨\underline{\overline{Keypad 0-9}}^{⟩} Highlight a page number ⟨\underline{\overline{Keypad -}}^{⟩} Highlight previous page number ⟨\underline{\overline{Keypad +}}^{⟩} Highlight next page number

⟨

Keypad Enter

⟩

Jump to the highlighted page

⟨\underline{\overline{C-Enter}}^{⟩} Scroll columns up ⟨\underline{\overline{S-Enter}}^{⟩} Scroll columns up ⟨\underline{\overline{Enter}}^{⟩} Scroll columns down ⟨\underline{\overline{B}}^{⟩} Jump -1 pages ⟨\underline{\overline{F}}^{⟩} Jump 1 pages

5 Comments about the user interface

5.1 Scales

The Scale menu which allows to view the document at different sizes is divided into two parts. The first part shows a list of available scale bases, the second part, separated by a line, lists the relative scales which are applied with respect to the selected scale base.

By default two scales bases are available, the Natural size and the Pixel based base. When choosing the Pixel based scale base a relative scale of 1.0 causes one postscript point to correspond to one pixel on the screen.

When viewing a document at a relative scale of 1.0 using the Natural size base the page should appear in its real size, as if printed on paper. For the Natural size base to work properly gv has to know the correct size of the root window. Automatic detection of this size unfortunately only provides approximate results, therefore it is best if it is provided by the user. To do so the resource

   GV.screenSize:  <width> x <height>

should be added to the SCREEN_RESOURCES property of the screen the document is viewed on, with width and height describing the width and height of the root window in units of millimeters. For instance by using the command

   echo "GV.screenSize: 396 x 291" | xrdb -override -screen

a width of 346 mm and a height of 291 mm will be used for the Natural size scale base. Alternatively the resource

   GV.screenSize_<machine>_<disp>_<scr>:  <width> x <height>

may be added to a resource file read by gv. Here machine ,disp and ANGLED describe the display on which gv displays the document. For instance, if the display is set to tic.tac.toe.wo:0.1 the resource should be specified as

   GV.screenSize_tic_0_1: 396 x 291

Note that this method doesn’t work on VMS. As a last alternative the resource

   GV.screenSize:  <width> x <height>

may be added to one of the resource files. However, for obvious reasons this method should be used only on single user machines.

5.2 Scrolling

It is probably a de facto standard for user interfaces that windows are supposed to be scrolled by means of scrollbars attached to their sides. However, for various reasons the use of scrollbars has been minimized in gv.

Instead, in all windows with obscured data, scrolling may be performed by

pressing button1, moving the mouse, then releasing button1

directly in the window. This includes the displayed page, zoom popups, the table of contents and the file and directory lists displayed in the file selection popup.

5.3 Displaying new versions of a document

During document creation it is usually desirable to always display the newest version of the document in work.

Instead of explicitly loading it via the file selection popup the most recent version can be displayed by pressing the Reload button in the main window. Choosing the Reload entry in the menu that pops up when clicking with the third mouse button anywhere on the displayed page certainly has the same effect.

More comfortable is the "Watch File" feature which may be switched on by selecting the corresponding entry in the State menu. If activated gv will check every now and then if a new version of the displayed file exists. If so it will be displayed automatically. By default the file is checked about once every second.

Finally it may also be left to the document creating program to trigger gv to update its display. To do so the program should send the SIGHUP signal to gv. For instance at the end of a shell script generating a postscript file from latex sources the line

   kill -SIGHUP <gv_pid>

may be added (here gv pid is the process id of gv). Executing the script and thereby creating a new version of the document will then also cause the result to be displayed instantaneously. Please note that this feature is available only on X11 R6 based systems.

6 Frequently asked questions

6.1 What happened to the scroll bars?

They have been removed in order to make more room for the image. The displayed portion of the image may be moved by dragging the image or the paner (panel between the "Save Marked" and << >> buttons) with the mouse. Alternatively, the cursor arrow keys can be used.

6.2 How can I make even more room for the image?

Use the "spartan" style by starting gv with the command

         gv --spartan

This removes the "Open", "Print", "Save", "Reload", and page marking buttons (they are still available from the "File" and "Page" menus) and replaces them with the document attribute controls, which are normally along the top.

6.3 Small characters are not very clear

Antialiasing can improve the display of bitmapped fonts (eg. from TeX) when displayed on a colour or greyscale screen. The same technique is used by xdvi. Note that antialiasing requires at least Ghostscript version 4.x.

Antialiasing can be turned on from the "State" menu, and can be made default by saving

the setting in State | gv Options....

6.4 The output is not refreshed automatically

There are two methods that can be used to save the contents of the window when it’s not currently displayed: backing store and backing pixmap. Some X-servers seem to support only backing store (eg. VAXstations) and some only backing pixmap (eg. some X-terminals, including EWS).

In order to force gv to use one method or the other, use the State | Setup Options ... menu and toggle the "Backing Pixmap" button. When selected/highlighted (normally the default), gv will use backing pixmap; otherwise it will use backing store. Select "Apply" to use a new setting and "Save" to make it the new default.

7 gv internals

7.1 Interface with ghostscript

7.1.1 Ghostview interface to ghostscript

When the GHOSTVIEW environment variable is set, ghostscript draws on an existing drawable rather than creating its own window. Ghostscript can be directed to draw on either a window or a pixmap.

7.1.2 Drawing on a Window

The GHOSTVIEW environment variable contains the window id of the target window. The window id is an integer. Ghostscript will use the attributes of the window to obtain the width, height, colormap, screen, and visual of the window. The remainder of the information is gotten from the GHOSTVIEW property on that window.

7.1.3 Drawing on a Pixmap

The GHOSTVIEW environment variable contains a window id and a pixmap id. They are integers separated by white space. Ghostscript will use the attributes of the window to obtain the colormap, screen, and visual to use. The width and height will be obtained from the pixmap. The remainder of the information, is gotten from the GHOSTVIEW property on the window. In this case, the property is deleted when read.

7.1.4 The GHOSTVIEW environment variable

parameters

          window-id [pixmap-id]

scanf format

          "%d %d"

Explanation of parameters:

window-id

tells ghostscript where to • read the GHOSTVIEW property

• send events

If pixmap-id is not present, ghostscript will draw on this window.

pixmap-id

If present, tells ghostscript that a pixmap will be used as the final destination for drawing. The window will not be touched for drawing purposes.

7.1.5 The GHOSTVIEW property

type STRING

parameters

          bpixmap orient llx lly urx ury xdpi ydpi [left bottom top right]

scanf format

          "%d %d %d %d %d %d %f %f %d %d %d %d"

Explanation of parameters:

pixmap pixmap id of the backing pixmap for the window. If no pixmap is to be used, this parameter should be zero. This parameter must be zero when drawing on a pixmap.

orient orientation of the page. The number represents clockwise rotation of the paper in degrees. Permitted values are 0, 90, 180, 270.

llx, lly, urx, ury

Bounding box of the drawable. The bounding box is specified in PostScript points in default user coordinates.

xdpi, ydpi

Resolution of window. (This can be derived from the other parameters, but not without roundoff error. These values are included to avoid this error.)

left, bottom, top, right (optional)

Margins around the window. The margins extend the imageable area beyond the boundaries of the window. This is primarily used for popup zoom windows. I have encountered several instances of PostScript programs that position themselves with respect to the imageable area. The margins are specified in PostScript points. If omitted, the margins are assumed to be 0.

7.1.6 Events from ghostscript

If the final destination is a pixmap, the client will get a property notify event when ghostscript reads the GHOSTVIEW property causing it to be deleted.

Ghostscript sends events to the window where it read the GHOSTVIEW property. These events are of type ClientMessage. The message type is set to either PAGE or DONE. The first long data value gives the window to be used to send replies to ghostscript. The second long data value gives the primary drawable. If rendering to a pixmap, it is the primary drawable. If rendering to a window, the backing pixmap is the primary drawable. If no backing pixmap is employed, then the window is the primary drawable. This field is necessary to distinguish multiple ghostscripts rendering to separate pixmaps where the GHOSTVIEW property was placed on the same window.

The PAGE message indicates that a "page" has completed. Ghostscript will wait until it receives a ClientMessage whose message type is NEXT before continuing.

The DONE message indicates that ghostscript has finished processing.

7.2 Comments recognized by ghostview

%!PS-Adobe-<real> [EPSF-<real>]
%%BoundingBox: <int> <int> <int> <int>|(atend)
%%CreationDate: <textline>
%%Orientation: Portrait|Landscape|(atend)
%%Pages: <uint>|(atend)
%%PageOrder: Ascend|Descend|Special|(atend)

%%Title: <textline>
%%DocumentMedia: <text> <real> <real> <real> <text> <text>
%%DocumentPageSizes: <text>
%%EndComments

Preview section:

%%BeginPreview
%%EndPreview

Defaults section:

%%BeginDefaults
%%PageBoundingBox: <int> <int> <int> <int>|(atend)
%%PageOrientation: Portrait|Landscape
%%PageMedia: <text>
%%EndDefaults

Prolog section:

%%BeginProlog
%%EndProlog

Setup section:

%%BeginSetup
%%PageBoundingBox: <int> <int> <int> <int>|(atend)
%%PageOrientation: Portrait|Landscape
%%PaperSize: <text>
%%EndSetup

Page properties:

%%Page: <text> <uint>
%%PageBoundingBox: <int> <int> <int> <int>|(atend)
%%PageOrientation: Portrait|Landscape
%%PageMedia: <text>
%%PaperSize: <text>
%%Trailer
%%EOF

Document section:

%%BeginDocument: <text> [<real>[<text>]]
%%EndDocument

Binary section:

%%BeginBinary: <uint>
%%EndBinary

Data section:

%%BeginData: <uint> [Hex|Binary|ASCII[Bytes|Lines]]
%%EndData

Chapter 7: gv internals 23

7.3 Paper Keywords and paper size in points

Letter           612x792
LetterSmall      612x792
Tabloid          792x1224
Ledger          1224x792
Legal            612x1008
Statement        396x612
Executive        540x720

A0               2384x3371
A1              1685x2384

A2              1190x1684
A3               842x1190
A4               595x842
A4Small          595x842
A5               420x595
B4               729x1032
B5               516x729
Envelope         ???x???
Folio            612x936
Quarto           610x780
10x14            720x1008

Appendix A GNU General Public License

Version 3, 29 June 2007

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007 Free Software Foundation, Inc. http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The GNU General Public License is a free, copyleft license for software and other kinds of works.

The licenses for most software and other practical works are designed to take away your freedom to share and change the works. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change all versions of a program—to make sure it remains free software for all its users. We, the Free Software Foundation, use the GNU General Public License for most of our software; it applies also to any other work released this way by its authors. You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for them if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs, and that you know you can do these things.

To protect your rights, we need to prevent others from denying you these rights or asking you to surrender the rights. Therefore, you have certain responsibilities if you distribute copies of the software, or if you modify it: responsibilities to respect the freedom of others.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must pass on to the recipients the same freedoms that you received. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

Developers that use the GNU GPL protect your rights with two steps: (1) assert copyright on the software, and (2) offer you this License giving you legal permission to copy, distribute and/or modify it.

For the developers’ and authors’ protection, the GPL clearly explains that there is no warranty for this free software. For both users’ and authors’ sake, the GPL requires that modified versions be marked as changed, so that their problems will not be attributed erroneously to authors of previous versions.

Some devices are designed to deny users access to install or run modified versions of the software inside them, although the manufacturer can do so. This is fundamentally incompatible with the aim of protecting users’ freedom to change the software. The systematic pattern of such abuse occurs in the area of products for individuals to use, which is precisely where it is most unacceptable. Therefore, we have designed this version of the GPL to prohibit the practice for those products. If such problems arise substantially in other domains, we stand ready to extend this provision to those domains in future versions of the GPL, as needed to protect the freedom of users.

Finally, every program is threatened constantly by software patents. States should not allow patents to restrict development and use of software on general-purpose computers, but in those that do, we wish to avoid the special danger that patents applied to a free program could make it effectively proprietary. To prevent this, the GPL assures that patents cannot be used to render the program non-free.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS

0. Definitions. “This License” refers to version 3 of the GNU General Public License. “Copyright” also means copyright-like laws that apply to other kinds of works, such as semiconductor masks. “The Program” refers to any copyrightable work licensed under this License. Each licensee is addressed as “you”. “Licensees” and “recipients” may be individuals or organizations. To “modify” a work means to copy from or adapt all or part of the work in a fashion requiring copyright permission, other than the making of an exact copy. The resulting work is called a “modified version” of the earlier work or a work “based on” the earlier work. A “covered work” means either the unmodified Program or a work based on the Program. To “propagate” a work means to do anything with it that, without permission, would make you directly or secondarily liable for infringement under applicable copyright law, except executing it on a computer or modifying a private copy. Propagation includes copying, distribution (with or without modification), making available to the public, and in some countries other activities as well. To “convey” a work means any kind of propagation that enables other parties to make or receive copies. Mere interaction with a user through a computer network, with no transfer of a copy, is not conveying. An interactive user interface displays “Appropriate Legal Notices” to the extent that it includes a convenient and prominently visible feature that (1) displays an appropriate copyright notice, and (2) tells the user that there is no warranty for the work (except to the extent that warranties are provided), that licensees may convey the work under this License, and how to view a copy of this License. If the interface presents a list of user commands or options, such as a menu, a prominent item in the list meets this criterion.

1. Source Code. The “source code” for a work means the preferred form of the work for making modifications to it. “Object code” means any non-source form of a work. A “Standard Interface” means an interface that either is an official standard defined by a recognized standards body, or, in the case of interfaces specified for a particular programming language, one that is widely used among developers working in that language.

The “System Libraries” of an executable work include anything, other than the work as a whole, that (a) is included in the normal form of packaging a Major Component, but which is not part of that Major Component, and (b) serves only to enable use of the work with that Major Component, or to implement a Standard Interface for which an implementation is available to the public in source code form. A “Major Component”, in this context, means a major essential component (kernel, window system, and so on) of the specific operating system (if any) on which the executable work runs, or a compiler used to produce the work, or an object code interpreter used to run it.

The “Corresponding Source” for a work in object code form means all the source code needed to generate, install, and (for an executable work) run the object code and to modify the work, including scripts to control those activities. However, it does not include the work’s System Libraries, or general-purpose tools or generally available free programs which are used unmodified in performing those activities but which are not part of the work. For example, Corresponding Source includes interface definition files associated with source files for the work, and the source code for shared libraries and dynamically linked subprograms that the work is specifically designed to require, such as by intimate data communication or control flow between those subprograms and other parts of the work.

The Corresponding Source need not include anything that users can regenerate automatically from other parts of the Corresponding Source.

The Corresponding Source for a work in source code form is that same work.

2. Basic Permissions.

All rights granted under this License are granted for the term of copyright on the Program, and are irrevocable provided the stated conditions are met. This License explicitly affirms your unlimited permission to run the unmodified Program. The output from running a covered work is covered by this License only if the output, given its content, constitutes a covered work. This License acknowledges your rights of fair use or other equivalent, as provided by copyright law.

You may make, run and propagate covered works that you do not convey, without conditions so long as your license otherwise remains in force. You may convey covered works to others for the sole purpose of having them make modifications exclusively for you, or provide you with facilities for running those works, provided that you comply with the terms of this License in conveying all material for which you do not control copyright. Those thus making or running the covered works for you must do so exclusively on your behalf, under your direction and control, on terms that prohibit them from making any copies of your copyrighted material outside their relationship with you.

Conveying under any other circumstances is permitted solely under the conditions stated below. Sublicensing is not allowed; section 10 makes it unnecessary.

3. Protecting Users’ Legal Rights From Anti-Circumvention Law.

No covered work shall be deemed part of an effective technological measure under any applicable law fulfilling obligations under article 11 of the WIPO copyright treaty adopted on 20 December 1996, or similar laws prohibiting or restricting circumvention of such measures.

When you convey a covered work, you waive any legal power to forbid circumvention of technological measures to the extent such circumvention is effected by exercising rights under this License with respect to the covered work, and you disclaim any intention to limit operation or modification of the work as a means of enforcing, against the work’s users, your or third parties’ legal rights to forbid circumvention of technological measures.

4. Conveying Verbatim Copies.

You may convey verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice; keep intact all notices stating that this License and any non-permissive terms added in accord with section 7 apply to the code; keep intact all notices of the absence of any warranty; and give all recipients a copy of this License along with the Program. You may charge any price or no price for each copy that you convey, and you may offer support or warranty protection for a fee.

5. Conveying Modified Source Versions. You may convey a work based on the Program, or the modifications to produce it from the Program, in the form of source code under the terms of section 4, provided that you also meet all of these conditions: a. The work must carry prominent notices stating that you modified it, and giving a

relevant date. b. The work must carry prominent notices stating that it is released under this Li-

cense and any conditions added under section 7. This requirement modifies the requirement in section 4 to “keep intact all notices”.

c. You must license the entire work, as a whole, under this License to anyone who comes into possession of a copy. This License will therefore apply, along with any applicable section 7 additional terms, to the whole of the work, and all its parts, regardless of how they are packaged. This License gives no permission to license the work in any other way, but it does not invalidate such permission if you have separately received it.

d. If the work has interactive user interfaces, each must display Appropriate Legal Notices; however, if the Program has interactive interfaces that do not display Appropriate Legal Notices, your work need not make them do so.

A compilation of a covered work with other separate and independent works, which are not by their nature extensions of the covered work, and which are not combined with it such as to form a larger program, in or on a volume of a storage or distribution medium, is called an “aggregate” if the compilation and its resulting copyright are not used to limit the access or legal rights of the compilation’s users beyond what the individual works permit. Inclusion of a covered work in an aggregate does not cause this License to apply to the other parts of the aggregate.

6. Conveying Non-Source Forms.

You may convey a covered work in object code form under the terms of sections 4 and 5, provided that you also convey the machine-readable Corresponding Source under the terms of this License, in one of these ways: a. Convey the object code in, or embodied in, a physical product (including a phys-

ical distribution medium), accompanied by the Corresponding Source fixed on a durable physical medium customarily used for software interchange.

b. Convey the object code in, or embodied in, a physical product (including a physical distribution medium), accompanied by a written offer, valid for at least three years and valid for as long as you offer spare parts or customer support for that product model, to give anyone who possesses the object code either (1) a copy of the Corresponding Source for all the software in the product that is covered by this License, on a durable physical medium customarily used for software interchange, for a price no more than your reasonable cost of physically performing this conveying of source, or (2) access to copy the Corresponding Source from a network server at no charge.

c. Convey individual copies of the object code with a copy of the written offer to provide the Corresponding Source. This alternative is allowed only occasionally and noncommercially, and only if you received the object code with such an offer, in accord with subsection 6b.

d. Convey the object code by offering access from a designated place (gratis or for a charge), and offer equivalent access to the Corresponding Source in the same way through the same place at no further charge. You need not require recipients to copy the Corresponding Source along with the object code. If the place to copy the object code is a network server, the Corresponding Source may be on a different server (operated by you or a third party) that supports equivalent copying facilities, provided you maintain clear directions next to the object code saying where to find the Corresponding Source. Regardless of what server hosts the Corresponding Source, you remain obligated to ensure that it is available for as long as needed to satisfy these requirements.

e. Convey the object code using peer-to-peer transmission, provided you inform other peers where the object code and Corresponding Source of the work are being offered to the general public at no charge under subsection 6d.

A separable portion of the object code, whose source code is excluded from the Corresponding Source as a System Library, need not be included in conveying the object code work. A “User Product” is either (1) a “consumer product”, which means any tangible personal property which is normally used for personal, family, or household purposes, or (2) anything designed or sold for incorporation into a dwelling. In determining whether a product is a consumer product, doubtful cases shall be resolved in favor of coverage. For a particular product received by a particular user, “normally used” refers to a typical or common use of that class of product, regardless of the status of the particular user or of the way in which the particular user actually uses, or expects or is expected to use, the product. A product is a consumer product regardless of whether the product has substantial commercial, industrial or non-consumer uses, unless such uses represent the only significant mode of use of the product. “Installation Information” for a User Product means any methods, procedures, authorization keys, or other information required to install and execute modified versions of a covered work in that User Product from a modified version of its Corresponding Source.

The information must suffice to ensure that the continued functioning of the modified object code is in no case prevented or interfered with solely because modification has been made. If you convey an object code work under this section in, or with, or specifically for use in, a User Product, and the conveying occurs as part of a transaction in which the right of possession and use of the User Product is transferred to the recipient in perpetuity or for a fixed term (regardless of how the transaction is characterized), the Corresponding Source conveyed under this section must be accompanied by the Installation Information. But this requirement does not apply if neither you nor any third party retains the ability to install modified object code on the User Product (for example, the work has been installed in ROM). The requirement to provide Installation Information does not include a requirement to continue to provide support service, warranty, or updates for a work that has been modified or installed by the recipient, or for the User Product in which it has been modified or installed. Access to a network may be denied when the modification itself materially and adversely affects the operation of the network or violates the rules and protocols for communication across the network. Corresponding Source conveyed, and Installation Information provided, in accord with this section must be in a format that is publicly documented (and with an implementation available to the public in source code form), and must require no special password or key for unpacking, reading or copying.

7. Additional Terms. “Additional permissions” are terms that supplement the terms of this License by making exceptions from one or more of its conditions. Additional permissions that are applicable to the entire Program shall be treated as though they were included in this License, to the extent that they are valid under applicable law. If additional permissions apply only to part of the Program, that part may be used separately under those permissions, but the entire Program remains governed by this License without regard to the additional permissions. When you convey a copy of a covered work, you may at your option remove any additional permissions from that copy, or from any part of it. (Additional permissions may be written to require their own removal in certain cases when you modify the work.) You may place additional permissions on material, added by you to a covered work, for which you have or can give appropriate copyright permission. Notwithstanding any other provision of this License, for material you add to a covered work, you may (if authorized by the copyright holders of that material) supplement the terms of this License with terms: a. Disclaiming warranty or limiting liability differently from the terms of sections 15

and 16 of this License; or b. Requiring preservation of specified reasonable legal notices or author attributions

in that material or in the Appropriate Legal Notices displayed by works containing it; or

c. Prohibiting misrepresentation of the origin of that material, or requiring that modified versions of such material be marked in reasonable ways as different from the original version; or d. Limiting the use for publicity purposes of names of licensors or authors of the

material; or e. Declining to grant rights under trademark law for use of some trade names, trade-

marks, or service marks; or f. Requiring indemnification of licensors and authors of that material by anyone who

conveys the material (or modified versions of it) with contractual assumptions of liability to the recipient, for any liability that these contractual assumptions directly impose on those licensors and authors.

All other non-permissive additional terms are considered “further restrictions” within the meaning of section 10. If the Program as you received it, or any part of it, contains a notice stating that it is governed by this License along with a term that is a further restriction, you may remove that term. If a license document contains a further restriction but permits relicensing or conveying under this License, you may add to a covered work material governed by the terms of that license document, provided that the further restriction does not survive such relicensing or conveying. If you add terms to a covered work in accord with this section, you must place, in the relevant source files, a statement of the additional terms that apply to those files, or a notice indicating where to find the applicable terms. Additional terms, permissive or non-permissive, may be stated in the form of a separately written license, or stated as exceptions; the above requirements apply either way.

8. Termination.

You may not propagate or modify a covered work except as expressly provided under this License. Any attempt otherwise to propagate or modify it is void, and will automatically terminate your rights under this License (including any patent licenses granted under the third paragraph of section 11). However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, you do not qualify to receive new licenses for the same material under section 10.

9. Acceptance Not Required for Having Copies. You are not required to accept this License in order to receive or run a copy of the Program. Ancillary propagation of a covered work occurring solely as a consequence of using peer-to-peer transmission to receive a copy likewise does not require acceptance.

However, nothing other than this License grants you permission to propagate or modify any covered work. These actions infringe copyright if you do not accept this License. Therefore, by modifying or propagating a covered work, you indicate your acceptance of this License to do so.

10. Automatic Licensing of Downstream Recipients. Each time you convey a covered work, the recipient automatically receives a license from the original licensors, to run, modify and propagate that work, subject to this License. You are not responsible for enforcing compliance by third parties with this License. An “entity transaction” is a transaction transferring control of an organization, or substantially all assets of one, or subdividing an organization, or merging organizations. If propagation of a covered work results from an entity transaction, each party to that transaction who receives a copy of the work also receives whatever licenses to the work the party’s predecessor in interest had or could give under the previous paragraph, plus a right to possession of the Corresponding Source of the work from the predecessor in interest, if the predecessor has it or can get it with reasonable efforts. You may not impose any further restrictions on the exercise of the rights granted or affirmed under this License. For example, you may not impose a license fee, royalty, or other charge for exercise of rights granted under this License, and you may not initiate litigation (including a cross-claim or counterclaim in a lawsuit) alleging that any patent claim is infringed by making, using, selling, offering for sale, or importing the Program or any portion of it.

11. Patents.

A “contributor” is a copyright holder who authorizes use under this License of the Program or a work on which the Program is based. The work thus licensed is called the contributor’s “contributor version”. A contributor’s “essential patent claims” are all patent claims owned or controlled by the contributor, whether already acquired or hereafter acquired, that would be infringed by some manner, permitted by this License, of making, using, or selling its contributor version, but do not include claims that would be infringed only as a consequence of further modification of the contributor version. For purposes of this definition, “control” includes the right to grant patent sublicenses in a manner consistent with the requirements of this License. Each contributor grants you a non-exclusive, worldwide, royalty-free patent license under the contributor’s essential patent claims, to make, use, sell, offer for sale, import and otherwise run, modify and propagate the contents of its contributor version. In the following three paragraphs, a “patent license” is any express agreement or commitment, however denominated, not to enforce a patent (such as an express permission to practice a patent or covenant not to sue for patent infringement). To “grant” such a patent license to a party means to make such an agreement or commitment not to enforce a patent against the party. If you convey a covered work, knowingly relying on a patent license, and the Corresponding Source of the work is not available for anyone to copy, free of charge and under the terms of this License, through a publicly available network server or other readily accessible means, then you must either (1) cause the Corresponding Source to be so available, or (2) arrange to deprive yourself of the benefit of the patent license for this particular work, or (3) arrange, in a manner consistent with the requirements of this License, to extend the patent license to downstream recipients. “Knowingly relying” means you have actual knowledge that, but for the patent license, your conveying the covered work in a country, or your recipient’s use of the covered work in a country, would infringe one or more identifiable patents in that country that you have reason to believe are valid.

If, pursuant to or in connection with a single transaction or arrangement, you convey, or propagate by procuring conveyance of, a covered work, and grant a patent license to some of the parties receiving the covered work authorizing them to use, propagate, modify or convey a specific copy of the covered work, then the patent license you grant is automatically extended to all recipients of the covered work and works based on it.

A patent license is “discriminatory” if it does not include within the scope of its coverage, prohibits the exercise of, or is conditioned on the non-exercise of one or more of the rights that are specifically granted under this License. You may not convey a covered work if you are a party to an arrangement with a third party that is in the business of distributing software, under which you make payment to the third party based on the extent of your activity of conveying the work, and under which the third party grants, to any of the parties who would receive the covered work from you, a discriminatory patent license (a) in connection with copies of the covered work conveyed by you (or copies made from those copies), or (b) primarily for and in connection with specific products or compilations that contain the covered work, unless you entered into that arrangement, or that patent license was granted, prior to 28 March 2007.

Nothing in this License shall be construed as excluding or limiting any implied license or other defenses to infringement that may otherwise be available to you under applicable patent law.

12. No Surrender of Others’ Freedom.

If conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot convey a covered work so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not convey it at all. For example, if you agree to terms that obligate you to collect a royalty for further conveying from those to whom you convey the Program, the only way you could satisfy both those terms and this License would be to refrain entirely from conveying the Program.

13. Use with the GNU Affero General Public License.

Notwithstanding any other provision of this License, you have permission to link or combine any covered work with a work licensed under version 3 of the GNU Affero General Public License into a single combined work, and to convey the resulting work. The terms of this License will continue to apply to the part which is the covered work, but the special requirements of the GNU Affero General Public License, section 13, concerning interaction through a network will apply to the combination as such.

14. Revised Versions of this License.

The Free Software Foundation may publish revised and/or new versions of the GNU General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Program specifies that a certain numbered version of the GNU General Public License “or any later version” applies to it, you have the option of following the terms and conditions either of that numbered version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of the GNU General Public License, you may choose any version ever published by the Free Software Foundation.

If the Program specifies that a proxy can decide which future versions of the GNU General Public License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Program.

Later license versions may give you additional or different permissions. However, no additional obligations are imposed on any author or copyright holder as a result of your choosing to follow a later version.

15. Disclaimer of Warranty.

THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PER-MITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EX-PRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFEC-TIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

16. Limitation of Liability.

IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MODIFIES AND/OR CONVEYS THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, IN-CIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUS-TAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAM-AGES.

17. Interpretation of Sections 15 and 16.

If the disclaimer of warranty and limitation of liability provided above cannot be given local legal effect according to their terms, reviewing courts shall apply local law that most closely approximates an absolute waiver of all civil liability in connection with the Program, unless a warranty or assumption of liability accompanies a copy of the Program in return for a fee.

END OF TERMS AND CONDITIONS

How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively state the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

      one line to give the program's name and a brief idea of what it does.
      Copyright (C) year name of author
      This program is free software: you can redistribute it and/or modify
      it under the terms of the GNU General Public License as published by
      the Free Software Foundation, either version 3 of the License, or (at
      your option) any later version.
      This program is distributed in the hope that it will be useful, but
      WITHOUT ANY WARRANTY; without even the implied warranty of
      MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the GNU
      General Public License for more details.
      You should have received a copy of the GNU General Public License
      along with this program.  If not, see http://www.gnu.org/licenses/.

Also add information on how to contact you by electronic and paper mail.

If the program does terminal interaction, make it output a short notice like this when it starts in an interactive mode:

      program Copyright (C) year name of author
      This program comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
      This is free software, and you are welcome to redistribute it
      under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, your program’s commands might be different; for a GUI interface, you would use an “about box”.

You should also get your employer (if you work as a programmer) or school, if any, to sign a “copyright disclaimer” for the program, if necessary. For more information on this, and how to apply and follow the GNU GPL, see http://www.gnu.org/licenses/.

The GNU General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Lesser General Public License instead of this License. But first, please

read http://www.gnu.org/philosophy/why-not-lgpl.html.

(Index is nonexistent)